OpenStudy (anonymous):

A box is to be made from a rectangular sheet of cardboard 70cm by 150cm by cutting equal squares out of the four corners and bending four flaps to make the sides of the box. (The box has no top.) What is the largest possible volume of the box?

OpenStudy (anonymous):

it is 96234.08959 cc (approx)

OpenStudy (anonymous):

how did you come up with that? :) need help on the method, thanks

OpenStudy (anonymous):

if a is the value of each side of the box, 5a^2=70*150 a=45.82575695 so,a^3=96234.08959